# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v1

@@ Appendix G General Bounds on the Charge Transport Capacity
However, this bound on the operator norm is not very useful in practice, as it is looser than (or equivalent to, for $i_{0}=L/2)$ the more intuitive bound999 This bound can be obtained by noting that $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\frac{L-i_{0}}{2}\hat{I}\|_{\text{op}}=L-i_{0},$ (77) and $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\hat{N}+\frac{i_{0}}{2}\hat{I}\|_{\text{op}}=i_{0},$ (78) where $\hat{I}$ is the identity operator.
@@ H.1 Estimates of the Matrix Elements and the Frobenius and Operator norms in the Localized Phase
Hence, for the squared Frobenius norm of the charge transport capacity operator $\hat{M}_{i_{0}}^{(0)}$, which up to a factor of $1/4$ coincides with the standard squared Frobenius norm of the matrix $M_{i_{0}}^{(0)}$ (cf. Appendix B), we obtain
@@ Appendix I Operator Norm Data for $\hat{M}^{(0)}$
Figure 21: Analysis of the operator norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{\mathrm{op}}$, for the noninteracting Anderson Model with open boundary conditions with $i_{0}=\lfloor L/2\rfloor$. Our calculations are done for 10000 disorder realizations. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{\mathrm{op}}$. Each panel represents different disorder strengths at $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes between $L=20$ and $L=100$, in steps of 10. (e): $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}$ as a function of $L$. Each curve represents a different disorder strength. In the low disorder limit, the operator norm of $\hat{M}^{(0)}$ approaches the curve $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}\approx L/4$ while $L\ll\xi$, showing different scaling from that of the Frobenius norm, which was shown in the main text in Fig. 9.
Figure 22: Analysis of the operator norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{\mathrm{op}}$, for the noninteracting quasiperiodic model with open boundary conditions over 500 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$. Here, $i_{0}=\lfloor L/2\rfloor$ and $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{\mathrm{op}}$. Each panel shows distributions at different potential depths, at $h=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size. We calculate distributions between $L=20$ and $L=100$ in steps of 10. (e): $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}$ as a function of $L$. Each curve represents a different potential depth $h$, ranging from $h=0.1$ to $h=3$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow).
@@
Figure 23: (a)-(d): Probability density functions of $\|\hat{M}^{(1)}\|$ in the interacting quasiperiodic model with $k=\sqrt{3}$. Within each panel, each distribution represents a different system size $L$. (e): Scaling of $\mathrm{med}\|\hat{M}^{(1)}\|$ with $L$. Each curve represents a different potential depth $h$. One sees that the spikes occur at different locations compared to the case $k=\sqrt{2}$ considered in the main text, notably at $L=40$ and 70. Those system sizes are approximately commensurate with the underlying potential, as $\sqrt{3}\times 41\approx 71$ and $\sqrt{3}\times 71\approx 123$.

### v2

@@ Appendix G General Bounds on the Charge Transport Capacity
However, this bound on the operator norm is not very useful in practice, as it is looser than (or equivalent to, for $i_{0}=L/2)$ the more intuitive bound101010 This bound can be obtained by noting that $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\frac{L-i_{0}}{2}\hat{I}\|_{\text{op}}=L-i_{0},$ (77) and $\|\Delta\hat{N}_{i_{0}}(t)\|_{\text{op}}\leq 2\|\hat{N}_{R}-\hat{N}+\frac{i_{0}}{2}\hat{I}\|_{\text{op}}=i_{0},$ (78) where $\hat{I}$ is the identity operator.
@@ H.1 Estimates of the Matrix Elements and the Frobenius and Operator norms in the Localized Phase
Hence, for the squared Frobenius norm of the charge transport capacity operator $\hat{M}_{i_{0}}^{(0)}$, which up to a factor of 1/41/41 / 4 coincides with the standard squared Frobenius norm of the matrix $M_{i_{0}}^{(0)}$ (cf. Appendix B), we obtain
@@ Appendix I Operator Norm Data for $\hat{M}^{(0)}$
Figure 21: Analysis of the operator norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{\mathrm{op}}$, for the noninteracting Anderson Model with open boundary conditions with $i_{0}=\lfloor L/2\rfloor$. The calculations here are done based on 10000 disorder realizations. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{\mathrm{op}}$. Each panel represents different disorder strengths at $W=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents different system sizes between $L=20$ and $L=100$, in steps of 10. (e): $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}$ as a function of $L$. Each curve represents a different disorder strength. In the low disorder limit, the operator norm of $\hat{M}^{(0)}$ approaches the curve $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}\approx L/4$ while $L\ll\xi$, showing different scaling from that of the Frobenius norm, which was shown in the main text in Fig. 9.
Figure 22: Analysis of the operator norm of the charge transport capacity, $\|\hat{M}^{(0)}\|_{\mathrm{op}}$, for the noninteracting quasiperiodic model with open boundary conditions based on 500 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=\frac{1}{k}$. Here, $i_{0}=\lfloor L/2\rfloor$ and $k=\sqrt{2}$. (a)–(d): Calculated probability density functions of $\|\hat{M}^{(0)}\|_{\mathrm{op}}$. Each panel shows distributions at different potential depths, at $h=0.6,1.6,2.6,3.1$. Within each panel, each colored curve represents a different system size. We calculate distributions between $L=20$ and $L=100$ in steps of 10. (e): $\mathrm{med}\|\hat{M}^{(0)}\|_{\mathrm{op}}$ as a function of $L$. Each curve represents a different potential depth $h$, ranging from $h=0.1$ to $h=3$ in steps of 0.5. We also show the curve at the transition $h=2$ (yellow).
@@
Figure 23: (a)-(d): Probability density functions of $\|\hat{M}^{(1)}\|$ in the interacting quasiperiodic model with $k=\sqrt{3}$, based on 500 evenly spaced phase shifts $\delta$ between $\delta=0$ and $\delta=1/k$. Within each panel, each distribution represents a different system size $L$. (e): Scaling of $\mathrm{med}\|\hat{M}^{(1)}\|$ with $L$. Each curve represents a different potential depth $h$. One sees that the spikes occur at different locations compared to the case $k=\sqrt{2}$ considered in the main text, notably at $L=40$ and 70. Those system sizes are approximately commensurate with the underlying potential, as $\sqrt{3}\times 41\approx 71$ and $\sqrt{3}\times 71\approx 123$.